Consider the closed tank system, consisting of three brine tanks. Tank 1 has volume 20 gallons, Tank 2 has volume 10 gallons, and Tank 3 has volume 20 gallons. Brine flows from Tank 1 to Tank 2, from Tank 2 to Tank 3, and from Tank 3 back to Tank 1 at the rate of 20 gallons per minute.

  1. a) Letting x1(t), x2(t), x3(t) be the amount of salt in each tank at time t, what is the first order system that x′1, x′2, x′3 satisfies?

  2. b) By finding the eigenvalues and eigenvectors of an appropriate matrix, write down the general solution of the first order system in part a).

  3. c) Initially there is 20 pounds of salt in Tank 1, and no salt in the other tanks. What is the limiting amount of salt in each tank as t → ∞?

could someone help with this homework question? I got stuck on it for like 3 hours :((( ty for any piece of help!
Answer #1

(2) Consider the closed tank systems, consisting of aree brine tanks. with volumes V₂, V2 and V3. The inflow to tank I is theHere , v =X3=20 gallows, z togallows and r=20gal 2-to gallons an it 20 10 z = 22013- 13 V V3 X2 = fx * x2 - 20 *4-20 Xz= X1-2² -3 -4x²5x = -x (x²+40+5)=0 2_ - [+(2+i) ] [ )+(2-1)]=0 de 20 a 3-2-1 /2 = -2+1 s engenvan For every i we find to own vectorgeneral son : Plt) = Ciel) + C2 Walt) +3Wz(t) Drit) = (20 tear-lo cost 2oite [ -62 cost-₂ cost) + (6 +asmet citezt (-62-63)times and in I tank -t and thank-3 in tank-2 - (couch)